# HeadGaS: Real-Time Animatable Head Avatars via 3D Gaussian Splatting

### v1

@@ 6 Conclusion
We presented HeadGaS, a model for animatable head reconstruction and rendering from a monocular video that renders on real-time. Our extensive evaluation showed that the proposed model results in state-of-the-art performance, surpassing the baselines, while rendering on real-time frame rates (over 200fps for $512^{2}$ resolution). We justified our design choices via a set of ablations, where we demonstrated that linearly blending implicit features leads to less artifacts than the alternative of blending explicit parameters directly. Moreover, we have shown that changing colors and opacity is more effective than the intuitive alternative of deforming the Gaussian mean positions.
@@
HeadGaS: Real-Time Animatable Head Avatars via 3D Gaussian Splatting
This supplementary material provides further details about HeadGaS, as well as additional results. Section A provides some insights on the learned 3D Gaussians, such as visualization of the learned feature basis as well as a gradual removal of the Gaussians to observe the occluded content. Section B provides more implementation details on the methods used in the ablation of the main paper. Further, we provide some time analysis in Section C. In Section D we show more qualitative results on the novel expression task comparing ours against baselines. Finally, in this supplement we refer the reader to a demo video which contains method highlights and a few result sequences.
@@ A Understanding the learned 3D Gaussians
Basis visualization
Since HeadGaS relies on a feature basis for blending, it would be beneficial to understand what this basis is learning. For this purpose, we utilize expression parameters as one-hot vectors to our model and show the generated image alongside the 3DMM mesh corresponding to that expression in Figure S1. The top section of the figure shows two examples on the NBS data [11], using Face Warehouse [5], while the bottom part provides two FLAME [20] based examples from the INSTA data [48]. Note that for the FLAME-based tracking we noticed that the optimized neutral expressions were far from a vector of zeros, therefore we normalized the one-hot expression vectors first, by adding the expression weights of a neutral expression from the training set. Since the Face Warehouse basis is more semantic, i.e. every expression element corresponds to a more local and interpretable action, such as winking, or opening mouth in surprise, we observe more drastic changes in this base compared to FLAME. The results show that HeadGaS learns a reasonable feature basis that aligns well with the 3DMM expressions. However, this is limited by the level in which an expression is observed in the training data, e.g. for subject 2 from the top we observe that winking does not work quite well (both eyes are closing instead of one) due to this reason.
Figure S1: Learned feature basis visualization. Top: Expression parameters from the Face Warehouse model Bottom: Expression parameters from the FLAME model.
Gradual removal of Gaussians
The purpose of this experiment is to reveal what the intermediate Gaussians in a particular frame represent. This is of interest because the proposed method relies on over-representation, i.e. multiple Gaussians will represent certain face areas (e.g. lips) and they will occasionally become transparent to reveal other areas underneath (e.g. teeth) as needed. Therefore we remove the optimized 3D Gaussians gradually, using the camera view as direction, from near to far. Figure S2 shows the original rendering for a given frame as well as the rendering after we have removed the most frontal Gaussians in the mouth area. As expected, we start seeing teeth in the intermediate layers of visibility, as these structures have been observed from other frames (e.g. when the person was talking or laughing). Additionally, we observe background color (white) when dis-occluding the regions between the nose and the upper lip. This is due to the fact that there is no need to represent the structures underneath, as they are never observed by any frame.
@@ B Ablation details
Ours w/o blending
This model aims to verify that using the expression parameters as a weight for blending Gaussian features works better than using it as a simple condition to the MLP. Therefore, here we do not have a basis of latent features $\bm{F}$. Instead, the expression vector $\bm{e}_{i}$ and the encoded position $\bm{\mu}$ are fed directly into $\phi(\cdot)$ to predict the color and opacity. The implementation of the MLP remains the same.
Ours w/o $\mathcal{L}_{p}$
This model is the same as the proposed HeadGaS and simply has the perceptual loss disabled.
@@ C Time analysis
Time vs performance
In Figure S3 we visualize the rendering speed in FPS vs image quality (PSNR) for HeadGaS and other baselines on the INSTA data [48]. The radius circles are proportional to training time (smaller is better). Note that the x-axis (FPS) is presented in logarithmic scale. We observe that the proposed method clearly outperforms others in both rendering speed and quality, while the training time is closer to the fastest methods (INSTA).
Figure S3: Rendering speed (FPS in logarithmic scale) vs PSNR plot comparing different methods on the INSTA dataset. The circle radius represents training time.
Rendering time vs image size
@@
Figure S4: Rendering time with respect to image resolution for HeadGaS on the INSTA dataset. We show the statistics of all subjects, with number of Gaussians ranging from 21k to 37k.
Ablation on number of 3D Gaussians and rendering time
@@ D Additional qualitative results
We provide additional qualitative results comparing HeadGaS against the most recent baselines [48, 47, 11] in Figure S5 and Figure S6. We observe that generally our model renders images with less artifacts, higher similarity to the ground truth expression, more noticeable reflecting glasses and skin specularities (Figure S5, row 2).
@@
Figure S6: Additional qualitative results on the INSTA data. From left to right: Ground truth, HeadGaS (ours), INSTA and PointAvatar.

### v2

@@ 6 Conclusion
We presented HeadGaS, a model for animatable head reconstruction and rendering from a monocular video that renders on real-time. Our extensive evaluation showed that the proposed model results in state-of-the-art performance, clearly surpassing the baselines, while rendering on real-time frame rates (about 250 fps for a $512^{2}$ resolution). We justified our design choices via a set of ablations, where we demonstrated that linearly blending implicit features leads to less artifacts than the alternative of blending explicit parameters directly. Moreover, we have shown that changing colors and opacity is more effective than the intuitive alternative of transforming the Gaussian mean positions. Future work can be dedicated to improving the memory efficiency of the HeadGaS feature bases.
@@
This supplementary material provides further details about HeadGaS, as well as additional results. Section A provides some insights on the learned 3D Gaussians, such as visualization of the learned feature basis as well as a gradual removal of the Gaussians to observe the occluded content. Section B provides more implementation details on the methods used in the ablation of the main paper. Further, we provide some time analysis in Section C. In Section D we show more qualitative results on the novel expression task comparing ours against baselines. Finally, in this supplement we kindly refer the reader to a demo video which contains method highlights and various result sequences as comparison with state-of-the-art methods, ablation and novel view synthesis.
@@ A Understanding the Learned 3D Gaussians
A.0.1 Basis visualization
Since HeadGaS relies on a feature basis for blending, it would be beneficial to understand what this basis is learning. For this purpose, we utilize expression parameters as one-hot vectors to our model and show the generated image alongside the 3DMM mesh corresponding to that expression in Figure S1. The left section of the figure shows two examples on the NBS data [14], using Face Warehouse [6], while the right part provides two FLAME [24] based examples from the INSTA data [62]. Note that for the FLAME-based tracking we noticed that the optimized neutral expressions were far from a vector of zeros, therefore we normalized the one-hot expression vectors first, by adding the expression weights of a neutral expression from the training set. Since the Face Warehouse basis is more semantic, i.e. every expression element corresponds to a more local and interpretable action, such as winking, or opening mouth in surprise, we observe more drastic changes in this base compared to FLAME. The results show that HeadGaS learns a reasonable feature basis that aligns well with the 3DMM expressions. However, this is limited by the level in which an expression is observed in the training data, e.g. for subject 2 from the top we observe that winking does not work quite well (both eyes are closing instead of one) due to this reason.
Figure S1: Learned feature basis visualization. Left: Expression parameters from the Face Warehouse model Right: Expression parameters from the FLAME model.
A.0.2 Gradual removal of Gaussians
The purpose of this experiment is to reveal what the intermediate Gaussians in a particular frame represent. This is of interest because the proposed method relies on over-representation, i.e. multiple Gaussians will represent certain face areas (e.g. lips) and they will occasionally become transparent to reveal other areas underneath (e.g. teeth) as needed. Therefore we remove the optimized 3D Gaussians gradually, using the camera view as direction, from near to far. Figure S2 shows the original rendering for a given frame as well as the rendering after we have removed the most frontal Gaussians in the mouth area. As expected, we start seeing teeth in the intermediate layers of visibility, as these structures have been observed from other frames (e.g. when the person was talking or laughing). This reflects that, our model does not simply re-color the Gaussians in those areas to accommodate teeth instead of lips, but rather has a separate set of Gaussians to represent the teeth. Additionally, we observe background color (white) when dis-occluding the regions between the nose and the upper lip. This is due to the fact that there is no need to represent the structures underneath, as they are never observed by any frame.
@@ A.0.2 Gradual removal of Gaussians
A.0.3 Effect of going outside of the training manifold
[…]
Figure S3: (Top) Navigation of 3DMM expression space for 1 parameter. (Bottom) Navigation of viewing angle.
@@ B.0.2 Ours w/ $\Delta(\mu,R)$
To facilitate convergence, we additionally add a regularization term to the estimated position shift, encouraging it to remain in a small range. The final loss then becomes
$\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{1}(I_{\text{r}},I_{\text{gt}})+\lambda_{s}\mathcal{L}_{\text{SSIM}}(I_{\text{r}},I_{\text{gt}})+\lambda_{p}\mathcal{L}_{p}(I_{\text{r}},I_{\text{gt}})+\lambda_{\mu}\mathcal{L}_{1}(\Delta\bm{\mu}).$ (10)
B.0.3 Ours change all
This model uses the learned feature basis $\bm{F}$ of HeadGaS to predict color, opacity as well as a shift of positions $\bm{\mu}$ and transformation of rotations $\bm{R}$. Thereby we feed the blended feature $\bm{f}_{i}$ to an MLP $\phi^{\prime\prime}(\cdot)$, which outputs sh colors, opacity, 3 values of position shift and 4 values of rotation (represented as quaternion). The transformations are applied in the same way as in Ours w/ $\Delta(\mu,R)$. Also here we apply the regularization term in the loss function, as in eq. (10).
B.0.4 Ours w/o blending
This model aims to verify that using the expression parameters as a weight for blending Gaussian features works better than using it as a simple condition to the MLP. Therefore, here we do not have a basis of latent features $\bm{F}$. Instead, the expression vector $\bm{e}_{i}$ and the encoded position $\bm{\mu}$ are fed directly into $\phi(\cdot)$ to predict the color and opacity. We hypothesize that this baseline requires more capacity for the MLP, as, in contrast to our proposed method, it has to learn all dynamics of the face at once. Therefore, we do not restrict our experiments to a small MLP of two layers, but rather extend its capacity until it reaches a plateau. Thus, the MLP here results in five linear layers, each followed by a leaky ReLU.
B.0.5 Ours w/o $\mathcal{L}_{p}$
This model is the same as the proposed HeadGaS and simply has the perceptual loss disabled
$\mathcal{L}_{\text{total}}=\lambda_{1}\mathcal{L}_{1}(I_{\text{r}},I_{\text{gt}})+\lambda_{s}\mathcal{L}_{\text{SSIM}}(I_{\text{r}},I_{\text{gt}}).$ (11)
@@ C Time analysis
C.0.1 Rendering time vs image size
@@
Figure S4: Rendering time versus resolution for HeadGaS on the INSTA dataset. We show the statistics of all subjects, with number of Gaussians ranging from 21k to 37k.
C.0.2 Ablation on number of 3D Gaussians and rendering time
@@ D Additional qualitative results
We provide additional qualitative results comparing HeadGaS against the most recent baselines [62, 61, 14, 4] in Figure S5 and Figure S6. We observe that generally our model renders images with less artifacts, higher similarity to the ground truth expression, more noticeable reflecting glasses and skin specularities (Figure S5, row 2).
@@
Figure S6: Additional qualitative results on the INSTA data. From left to right: Ground truth, HeadGaS (ours), INSTA, PointAvatar and FLARE.